OpenStudy (anonymous):

What is the sale price of an item that is 10% off $14.50? $1.45 $13.05 $14.40 $15.95

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

smh

OpenStudy (harsimran_hs4):

step 1 : calculate 10% of 14.50 10 % of 14.50 = 10/100 * 14.50 = 1.45 step 2: SP = Cost - discount SP = 14.50 - 1.45 = 13.05

OpenStudy (anonymous):

Change a percent to a decimal, move the decimal point two places to the left and drop the percent sign $14.50 * .1 = $1.45. $14.50 - $1.45 = 13.05.